# Prompting as Scientific Inquiry

### v1

@@ 1 Introduction
Wei et al. (2022) showed that using in-context learning to generate human-like reasoning chains resulted in higher math scores for LLMs, while Kojima et al. (2022) built on the tradition of exploratory prompting popularized by GPT-2’s prompting studies (e.g., using "TL;DR" for generating summaries (Radford et al., 2019)), to show this could be done by simply appending five words to a prompt: “let’s think step-by-step.” The intuition was that strings used to signal when humans would generate reasoning chains could prompt models to enact this behavior. The key insight wasn’t that models simply needed more raw computational power or an method to engineer such computation into a model manually, though later studies showed that models could use inference-time compute without “thinking” out loud (Pfau et al., 2024). Researchers instead took advantage of behavioral patterns models learned from their training corpora, leveraging the correlations between reasoning chains and correct outputs. This discovery has since spawned a new generation of training procedures that leverage inference-time compute strategies. Modern systems like DeepSeek-R1 (Guo et al., 2025) and GPT-4o1 (OpenAI, 2024) now demonstrate that models can reason better in recognizable ways—achieving improved results on complex mathematical problems and challenging domain-specific tasks such as graduate-level questions in GPQA (Rein et al., ). Asking models to show their work revealed how we could train models to do something more like thinking, a process of exploration and refinement that would have been difficult to achieve without prompting.
@@ 3 How did Prompting Get Counted Out?
First, a series of studies demonstrated the sensitivity of prompting (Wallace et al., 2019; Arora et al., 2022; Min et al., 2022, inter alia). Model behavior can be unpredictable, counterintuitive, and small changes can change the performance of a system drastically. As a result, researcher assumed that prompting cannot be trustworthy or useful, and many suggested we need to find a way to get rid of it (Meincke et al., 2025; Kosch & Feger, 2025; Morris, 2024, inter alia).
@@ 4 Prompting as Scientific Inquiry
Putting LLMs in new contexts, examining the generated outcomes, and attempting to find patterns in the resulting probability distributions, these are precisely the necessary ingredients of exploratory science. And exploratory science is vital when we are confronted with a new system that evades intuitive explanation, as we are today. A similar point was made eight years ago in Neuroscience, by a paper titled “Could a neuroscientist understand a microprocessor?” (Jonas & Kording, 2017). The answer is “No.” precisely because the strategies taken would not exploit the regularity with which these systems could be manipulated. LLMs have been handed to us on a silver platter, from this perspective: they already speak approximately the same language we do, shouldn’t we be excited to be able to probe them with it?
@@ 5.2 Practical Implications
Scientific Narrative. Many key discoveries in interpretability begin with behavioral anomalies. For instance, the characterization of induction heads followed observations of copying behavior in Transformers (Al-Rfou et al., 2019; Elhage et al., 2021). Prompting offers a structured hypothesis space to surface such behaviors, guiding mechanistic follow-up. In this light, prompting is not just an evaluation tool; it drives scientific discovery by revealing what is worth interpreting.

### v2

@@ 1 Introduction
Wei et al. (2022) showed that using in-context learning to generate human-like reasoning chains resulted in higher math scores for LLMs, while Kojima et al. (2022) built on the tradition of exploratory prompting popularized by GPT-2’s prompting studies (e.g., using "TL;DR" for generating summaries (Radford et al., 2019)), to show this could be done by simply appending five words to a prompt: “let’s think step-by-step.” The intuition was that strings used to signal when humans would generate reasoning chains could prompt models to enact this behavior. The key insight wasn’t that models simply needed more raw computational power or an method to engineer such computation into a model manually, though later studies showed that models could use inference-time compute without “thinking” out loud (Pfau et al., 2024). Researchers instead took advantage of behavioral patterns models learned from their training corpora, leveraging the correlations between reasoning chains and correct outputs. This discovery has since spawned a new generation of training procedures that leverage inference-time compute strategies. Modern systems like DeepSeek-R1 (Guo et al., 2025) and GPT-4o1 (OpenAI, 2024) now demonstrate that models can reason better in recognizable ways—achieving improved results on complex mathematical problems and challenging domain-specific tasks such as graduate-level questions in GPQA (Rein et al., 2024). Asking models to show their work revealed how we could train models to do something more like thinking, a process of exploration and refinement that would have been difficult to achieve without prompting.
@@ 3 How did Prompting Get Counted Out?
First, a series of studies demonstrated the sensitivity of prompting (Wallace et al., 2019; Arora et al., 2022; Min et al., 2022, inter alia). Model behavior can be unpredictable, counterintuitive, and small changes can change the performance of a system drastically. As a result, researchers assumed that prompting cannot be trustworthy or useful, and many suggested we need to find a way to get rid of it by making LLMs more like traditional computational systems (Meincke et al., 2025; Kosch & Feger, 2025; Morris, 2024, inter alia).
@@ 4 Prompting as Scientific Inquiry
Putting LLMs in new contexts, examining the generated outcomes, and attempting to find patterns in the resulting probability distributions, these are precisely the necessary ingredients of exploratory science. This view is echoed by Shah (2025), who frames prompting as a legitimate form of behavioral science and urges the adoption of research norms from qualitative inquiry—such as systematic labeling, human-in-the-loop triangulation, and reproducibility protocols. Similarly, Liu et al. (2023) were among the first to formalize prompting as a structured method of engaging pretrained models, positioning it not as a workaround but as a new learning paradigm centered on input-output behavior.
Exploratory science is vital when we are confronted with a new system that evades intuitive explanation, as we are today. A similar point was made eight years ago in Neuroscience, by a paper titled “Could a neuroscientist understand a microprocessor?” (Jonas & Kording, 2017). The answer is “No.” precisely because the strategies taken would not exploit the regularity with which these systems could be manipulated. LLMs have been handed to us on a silver platter, from this perspective: they already speak approximately the same language we do, shouldn’t we be excited to be able to probe them with it?
@@ 5.2 Practical Implications
Scientific Narrative. Many key discoveries in interpretability begin with behavioral anomalies. For instance, the characterization of induction heads followed observations of copying behavior in Transformers (Al-Rfou et al., 2019; Elhage et al., 2021). Prompting offers a structured hypothesis space to surface such behaviors, guiding mechanistic follow-up. In this light, prompting is not just an evaluation tool; it drives scientific discovery by revealing what is worth interpreting. Shah (2025) proposes an analogy to qualitative social science: researchers intervene via language, document patterns, and triangulate model behavior through iterative experimentation. Such perspectives support our view that prompting reveals not just system capabilities but emergent regularities that merit formal study.